Shelly is shopping for laundry detergent,mustard,and canned tuna.She is trying to decide which of two products is the better buy. Using the following information, can you help Shelly?
Laundry detergent A $2.00 for 37 ounces
Laundry detergent B $2.37 for 38 ounces
Mustard A $.88 for 6 ounces
Mustard B $1.61 for 12 1/2 ounces
Canned tuna A $1.09 for 6 ounces
Canned tuna B $1.29 for 8 3/4 ounces

Answer by Theo(13342)

the most straight forward way is to find the cost per ounce.

detergent:
2.00 / 37 = .054...
2.37 / 38 = .062...
laundry detergent A is cheaper.

mustard:
.88 / 6 = .146...
1.61 / 12.5 = .1288...
mustard B is cheaper

canned tuna:
1.09/6 = .1816...
1.29/8.75 = .147...
canned tuna B is cheaper.

another way is to do a ratio analysis.

detergent:
if 37 ounces of detergent A costs 2.00, then 38 ounces of detergent A would cost 38/37 * 2.00 = 2.05...
since 38 ounces of detergent B cost 2.37, then detergent A is cheaper.
mustard:
if 6 ounces of mustard A cost .88, then 12.5 ounces of mustard A would cost 12.5/6 * .88 = 1.833...
since 12.5 ounces of mustard B cost 1.61, then mustard B is cheaper.

canned tuna:
if 6 ounces of canned tuna A cost 1.09, then 8.75 ounces of canned tuna A would cost 8.75 / 6 * 1.09 = 1.589...
since 8.75 ounces of canned tuna B cost 1.29, then canned tuna B is cheaper.

both methods will yield the same answer.